# Towards Immersive Mixed Reality Street Play

### v1

@@ Abstract.
We’re witnessing an upcoming paradigm shift as Mixed Reality (MR) See-through Head-Mounted Displays (HMDs) become ubiquitous, with use shifting from controlled, private settings to spontaneous, public ones. While location-based pervasive mobile games like Pokémon GO have seen success, the embodied interaction of MRHMDs is moving us from phone-based screen-touching gameplay to MRHMD-enabled collocated bodily play. Major tech companies are continuously releasing visionary videos where urban streets transform into vast mixed reality playgrounds—imagine Harry Potter-style wizard duels on city streets. However, few researchers have conducted real-world, in-the-wild studies of such Immersive Mixed Reality Street Play (IMRSP) in public spaces in anticipation of a near future with prevalent MRHMDs. Through empirical studies on a series of research-through-design game probes called Multiplayer Omnipresent Fighting Arena (MOFA), we gain initial understanding of this under-explored area by identifying the social implications, challenges, and opportunities of this new paradigm.
Social Augmented Reality, Collaborative Mixed Reality, Bystander Inclusion, Social Acceptance, Game Design, Urban Games, Pervasive Games, Collocated Bodily Play, See-through Head-mounted Display, In-the-wild research, Public Spaces, Spontaneous Encounters, Awkwardness, Safety Concerns
journalyear: 2024copyright: rightsretainedconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06conference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interaction design theory, concepts and paradigmsccs: Human-centered computing Mixed / augmented realityccs: Human-centered computing Collaborative and social computing systems and tools
Figure 1. A research-through-design game probe series, Multiplayer Omnipresent Fighting Arena (MOFA), for studying Immersive Mixed Reality Street Play. (a) “The Training” in the public park in Brooklyn; (b) “The Duel” on the urban streets of SoHo, Manhattan; (c) “The Dragon” in public atrium at the World Trade Center Transportation Hub of New York City.
@@ 1. Introduction
Major tech companies such as Apple and Meta are rapidly advancing Mixed Reality Head-Mounted Display (MRHMD) technology: Apple’s Vision Pro [^1] http://www.apple.com/newsroom/2023/06/introducing-apple-vision-pro/, visited on October 25, 2024 offers video see-through mixed reality in a headset form-factor, while Meta’s Orion [^2] https://about.fb.com/news/2024/09/introducing-orion-our-first-true-augmented-reality-glasses/, visited on October 25, 2024 boasts optical see-through capabilities in an eyeglass form. Meta CEO Mark Zuckerberg has even predicted that “smart glasses will replace phones by 2030” [^3] https://www.forbes.com/sites/davidbirch/2024/10/16/mark-zuckerberg-and-why-smart-glasses-will-replace-mobile-phones, visited on October 25, 2024. Some scholars forecast that augmented reality may become virtually indistinguishable from physical reality (Nijholt, 2023; Laato et al., 2021; Rauschnabel, 2021), underscoring the importance of understanding user experiences and social dynamics surrounding these emerging devices. As Itoh et al. (Itoh et al., 2021) demonstrate, passthrough technologies — whether video see-through or optical see-through— enable users to maintain awareness of their real-world surroundings and move freely, extending potential usage from controlled private settings to spontaneous public scenarios. This shift raises pressing questions about how ubiquitous MRHMD adoption could transform everyday life in public spaces in the near future.
In the current augmented reality landscape, Pokémon GO stands as the most successful application, generating \$6 billion over eight years (Iqbal, 2024). While promotional videos depict a future where people engage in device-free bodily play on public streets, the game in today’s market still primarily relies on traditional smartphone screen-touching interactions. Moreover, some scientific studies estimate that less than 10% of active Pokémon GO players use the game’s AR features (Laato et al., 2021). Despite this, Niantic, the company behind Pokémon GO, continues to release promotional videos  [^4] For one such video, see ”Lightship x Snapdragon Spaces — Niantic” https://www.youtube.com/watch?v=sDNjVxX0LrE, visited on October 25, 2024 showcasing a future where people engage in MRHMD-based bodily play on public streets, interacting directly with nearby friends 555”Niantic — Meet You Out There” https://www.youtube.com/watch?v=HljcLVXAxAU, visited on October 25, 2024. We term this vision ”Immersive Mixed Reality Street Play” (IMRSP)—a form of collocated bodily play in public spaces using MRHMDs that primarily involves body movements as the main mode of interaction. Although the technology to implement such games exists, there’s a significant research gap regarding empirical studies for MRHMD-based bodily play in the wild. Therefore, in this study, we explore the feasibility of this vision through the following research question (RQ):
[…]
@@
Gameplay is widely anticipated to be one of the “killer apps” of this envisioned MR future (Rhodes, 2019). Even so, the most successful Augmented Reality (AR) title to date—Pokémon GO, which has generated over \$6 billion in eight years (Iqbal, 2024)—still relies predominantly on smartphone-based, screen-touch interactions. In fact, scientific studies estimate that fewer than 10% of its active players make use of the game’s AR features (Laato et al., 2021). Nevertheless, Niantic, the game’s developer, continues to release speculative promotional videos666“Niantic — Meet You Out There”, https://www.youtube.com/watch?v=HljcLVXAxAU, visited on October 25, 2024 and tech demo videos777 “Lightship x Snapdragon Spaces — Niantic,” https://www.youtube.com/watch?v=sDNjVxX0LrE, visited on October 25, 2024 that depict a future in which people seamlessly engage in public, bodily gameplay using see-through MR devices—or even device-free interfaces—spontaneously interacting with friends in city streets. We refer to this scenario as Immersive Mixed Reality Street Play (IMRSP), a form of collocated bodily play in public spaces using MRHMDs that primarily involves body movements as the main mode of interaction. Although such visions have been popularized in media and the technology to implement these visions exists, real-world ”in-the-wild” deployments of MRHMD-based bodily play involving spontaneous users in city streets remain underexplored. Therefore, in this study, we explore the feasibility of this vision through the following research question:
@@
(RQ) What are the social implications if immersive mixed reality street play becomes a prevalent form of play in public spaces in an future with ubiquitous see-through head-mounted displays?
[…]
Our empirical studies provide an initial yet foundational understanding of Immersive Mixed Reality Street Play. By examining how MRHMD technologies reshape social norms, dynamics, interactions, and collaborative play in public spaces, our findings inform both future research directions and design guidelines for creating more responsible and inclusive public IMRSP experiences.
[…]
@@
Unlike traditional mobile-based AR games, IMRSP requires extensive physical body movement in public spaces. While this enables more immersive presence and collaborative excitement, it also introduces distinct ethical, acceptance, privacy, and safety concerns (Millard et al., 2024; Katins et al., 2024). The online review feedback revealed more concerns than those expressed by on-site participants. As IMRSP games are physically demanding and often challenge established social norms — due to the power imbalance between MRHMD wearers and bystanders (Chung et al., 2023) — the long-term player retention of IMRSP in public spaces remains uncertain. Those implications underscores that the road to ubiquitous IMRSP vision may involve friction unless we address these issues head-on.
@@
  • An initial real-world exploration of social dynamics and phenomena in IMRSP, revealing key benefits and challenges of this novel form of play.
[…]
    A comparative analysis of online reviews versus onsite gameplay experiences, highlighting the friction of social acceptability between idealized visions and practical reality.
  • Design suggestions and a research agenda for future IMRSP development derived from our exploratory work.
  • MOFA: An open-source spontaneous collocated multiplayer MR game framework to enable further research exploration for IMRSP.
2. Background And Related Works
@@ 2.1. Technology-mediated Street Play
Street play has long been an integral part of urban culture, transforming streets into dynamic spaces where social interactions, creativity, and spontaneity flourish (Tranter and Doyle, 1996). Traditional street play, characterized by informal, unstructured activities like hopscotch, tag, or ball games (Wridt, 2004), utilizes public spaces in ways that challenge the boundaries of social and spatial order within urban environments (Stevens, 2007). Streets become arenas for social interaction, where norms are negotiated and social bonds are strengthened (Whyte et al., 1980). The spontaneous nature of street play fosters inclusivity and accessibility, enabling people of various ages and backgrounds to participate and interact. Meanwhile, the fields of urban play (Chisik et al., 2022), playable city (Nijholt, 2016), pervasive play (Ahn et al., 2016), alternate reality games (Kim et al., 2008), and pervasive games (Montola et al., 2009) have all contributed to redefining human-city interactions  (Lee et al., 2021).
@@
Technological advances in network coverage, power consumption, and device miniaturization have paved the way for new forms of outdoor interactive computing (Balestrini et al., 2020). These breakthroughs have spawned diverse deployments and studies of technology use in public spaces. Examples include large interactive displays (Cao et al., 2008; O’Hara et al., 2008a), mobile-based mixed reality (Flintham et al., 2003), location-based gaming (such as ”Can You See Me Now?” (Benford et al., 2006), geocaching (O’Hara, 2008), and Pokémon GO (Paavilainen et al., 2017)), and human-robot interactions on streets (Pelikan et al., 2024). These technological interventions have transformed street play, introducing innovative forms of engagement that blend physical and digital realms seamlessly (De Souza e Silva, 2009).
[…]
Informed by such studies, HCI has developed conceptual frameworks for thinking about design for interactions in public, including research in the wild (Chamberlain et al., 2012), performance-led research approaches (Benford et al., 2013a), and design considerations of bystanders and spectators in public interactions (Reeves et al., 2005a; Reeves, 2011). These methodologies have been instrumental in understanding and shaping user experiences in outdoor settings.
[…]
@@
HCI research has documented numerous social phenomena that arise when people interact with technology in public contexts, such as public displays (O’Hara et al., 2008b), and public installation (Balestrini et al., 2016). Among these are the honeypot effect (Wouters et al., 2016), where an initial user attracts onlookers and encourages them to engage; showmanship and performance, where users become aware of an audience and amplify their actions  (Benford and Giannachi, 2011; Peltonen et al., 2008; Reeves et al., 2005b); social mirroring, in which newcomers imitate observed behaviors (Peltonen et al., 2008). Researchers have also noted social discomfort (Benford et al., 2012), social facilitation (improved performance in front of others) (Peltonen et al., 2008; Miller et al., [n. d.]), social awkwardness (Huggard et al., 2013; Press and Erel, 2023), and inhibition (anxiety or reluctance to participate publicly) (Miller et al., [n. d.]).
2.2. Technology-supported Collocated Bodily Play
@@
In recent years, HCI research has undergone a ”somatic turn” (Loke and Schiphorst, 2018). This shift places the human body at the center of interactive technology design, moving beyond traditional user-centered approaches (Dix, 2017) to focus on the body as the primary site of interaction and experience (Höök, 2018). Merleau-Ponty argues that the human body is fundamental to perception and action (Merleau-Ponty, 2008), with proprioception and kinesthesia serving as core movement-related senses (Zhou et al., 2021). Homewood et al. (Homewood et al., 2021) outline three evolving conceptions of the body in HCI: the individual body (embodied interaction), multiple bodies (plurality of experiences), and more-than-human bodies (entangled assemblages). These concepts, particularly entangled assemblages (Forlano, 2017), intercorporeality (Meyer et al., 2017), and post-phenomenology (Wakkary et al., 2018), provide theoretical frameworks for exploring technology-supported collocated bodily play.
@@
The conceptual frameworks of proxemics (Ciolek, 1983) and f-formations (Barua et al., 2021) significantly inform the practical exploration of collocated bodily play. They offer insights into how spatial and social arrangements enhance interaction within shared physical spaces. These principles are instrumental in designing collocated bodily experiences that are socially engaging and spatially harmonious. Research on collocated bodily play in exertion games (Kaos et al., 2019) further underscores the importance of multiplayer experiences and social interaction in enhancing game adherence and enjoyment. Collocated bodily play design is rooted in social bodily play (Mueller et al., 2018b, 2019), encompassing three key areas: bodily play (Mueller et al., 2014; Jørgensen and Kaygan, 2024; Martin-Niedecken, 2018; Mueller et al., 2020; Márquez Segura et al., 2013; Isbister et al., 2018), movement-based design (Mueller and Isbister, 2014; van Rheden et al., 2024; Van Delden et al., 2023; Garner et al., 2013; Hämäläinen et al., 2015), and social exertion games (Mueller et al., 2017; Sinclair et al., 2007; Márquez Segura et al., 2021; Mueller and Young, 2017; Elvitigala et al., 2024). This body of work collectively emphasizes the integration of physical activity, social interaction, and immersive digital content in creating engaging and harmonious social play experiences.
2.3. Collocated Mixed Reality

### v2

@@ Abstract.
As see-through Mixed Reality Head-Mounted Displays (MRHMDs) proliferate, their usage is gradually shifting from controlled, private settings to spontaneous, public contexts. While location-based augmented reality mobile games such as Pokémon GO have been successful, the embodied interaction afforded by MRHMDs moves play beyond phone-based screen-tapping toward co-located, bodily, movement-based play. In anticipation of widespread MRHMD adoption, major technology companies have teased concept videos envisioning urban streets as vast mixed reality playgrounds—imagine Harry Potter–style wizard duels in city streets—which we term Immersive Mixed Reality Street Play (IMRSP). However, few real-world studies examine such scenarios. Through empirical, in-the-wild studies of our research-through-design game probe, Multiplayer Omnipresent Fighting Arena (MOFA), deployed across diverse public venues, we offer initial insights into the social implications, challenges, opportunities, and design recommendations of IMRSP. The MOFA framework, which includes three gameplay modes—“The Training”, “The Duel”, and “The Dragon”—is open-sourced at https://github.com/realitydeslab/mofa.
Social Augmented Reality, Collaborative Mixed Reality, Co-located Mixed Reality, Bystander Inclusion, Social Acceptance, Game Design, Urban Games, Pervasive Games, Co-located Bodily Play, See-through Head-mounted Display, Research in the Wild, Public Spaces, Spontaneous Encounters, Awkwardness, Safety Concerns
copyright: ccjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: 7article: CSCW498publicationmonth: 11doi: 10.1145/3757679ccs: Human-centered computing Interaction design theory, concepts and paradigmsccs: Human-centered computing Mixed / augmented realityccs: Human-centered computing Collaborative and social computing systems and tools
Figure 1. Multiplayer Omnipresent Fighting Arena (MOFA) is a research-through-design game probe series studying the social implications of Immersive Mixed Reality Street Play (IMRSP)—the scenario of co-located bodily play with mixed reality see-through head-mounted displays in public spaces. We deployed MOFA in the wild across three public venues: (a) ”The Training” in a public park in Dumbo, Brooklyn; (b) ”The Duel” on the urban streets of SoHo, Manhattan; and (c) ”The Dragon” in the public atrium of World Trade Center Transportation Hub, New York City.
@@ 1. Introduction
Major technology companies such as Apple and Meta are rapidly advancing see-through Mixed Reality Head-Mounted Display (MRHMD) technology: Apple’s Vision Pro (Apple Inc., 2023) offers video see-through mixed reality in a headset form factor, while Meta’s Orion (Meta Platforms, Inc., 2024) boasts optical see-through capabilities in an eyeglasses‑style form factor. Meta CEO Mark Zuckerberg has even predicted that “smart glasses will replace phones by 2030” (Birch, 2024). Some scholars forecast that augmented reality could become virtually indistinguishable from physical reality in certain contexts (Nijholt, 2023; Laato et al., 2021; Rauschnabel, 2021), underscoring the importance of understanding user experiences and social dynamics surrounding these emerging devices. As Itoh et al. (Itoh et al., 2021) demonstrate, passthrough technologies—whether video see-through or optical see-through—enable users to maintain awareness of their real-world surroundings while moving freely, extending potential usage from controlled private settings to spontaneous public scenarios. This shift raises pressing questions about how ubiquitous MRHMD adoption could transform everyday public life in the near future.
@@
Games are widely anticipated to be one of the “killer apps” of this envisioned MR future (Rhodes, 2019). The most successful augmented reality title to date is Pokémon GO, which has generated over \$6 billion in eight years (Iqbal, 2024). While Pokémon GO’s promotional videos111“Discover Pokémon in the Real World with Pokémon GO!”, https://www.youtube.com/watch?v=2sj2iQyBTQs, visited on October 25, 2024 depict a future where people engage in device-free bodily play on public streets, games in today’s market still primarily rely on traditional smartphone screen-touching interactions. Moreover, some scientific studies estimate that less than 10% of active Pokémon GO players use the game’s AR features (Laato et al., 2021). Despite this, Niantic, the game’s developer, continues to release speculative promotional videos222“Niantic — Meet You Out There”, https://www.youtube.com/watch?v=HljcLVXAxAU, visited on October 25, 2024 and tech demo videos333 “Lightship x Snapdragon Spaces — Niantic,” https://www.youtube.com/watch?v=sDNjVxX0LrE, visited on October 25, 2024 that depict a future in which people seamlessly engage in bodily gameplay using see-through MR devices—or even device-free interfaces—spontaneously interacting with friends on city streets. We refer to this scenario as Immersive Mixed Reality Street Play (IMRSP), a form of co-located bodily play in public spaces using MRHMDs that primarily involves body movements as the main mode of interaction. IMRSP is not a new concept and has been popularized in mainstream media, including commercial apps such as Pokémon GO and science fiction films such as Free Guy. Although the MR technology to implement IMRSP exists, real-world deployments—particularly MRHMD-based bodily play involving spontaneous participants in urban environments—remain underexplored. Therefore, in this study, we explore this vision through the following research question (RQ):
@@
What social implications might arise when immersive mixed reality street play becomes prevalent in public spaces in a future with ubiquitous see‑through head‑mounted displays?
We developed Multiplayer Omnipresent Fighting Arena (MOFA) [^4] Fun fact: The pronunciation of “MOFA” coincidentally means “magic” in Chinese. as a research-through-design (Zimmerman et al., 2007) game probe to investigate IMRSP in real-world contexts. MOFA is an open-source [^5] MOFA is open-sourced at https://github.com/realitydeslab/mofa. MR game framework that goes beyond traditional location-based, screen-touching mobile gaming by enabling co-located, multiplayer bodily play through MRHMDs. To better support in-the-wild studies, we introduced MOFA’s “omnipresent” feature: unlike other co-located MR games that require local Wi-Fi, this feature enables gameplay to occur spontaneously anywhere—even without Wi-Fi, cellular, or Internet connectivity—by using Apple’s Multipeer Connectivity technology, as detailed in Section 3. The framework centers on spell-casting gestures with a physical wand. This magic-themed interaction serves three key purposes: First, the spell-casting gestures become intuitive once users recognize the Harry Potter–inspired context, even if the movements initially seem unusual; Second, ranged attacks using virtual magical energy balls avoid physical contact, improving safety in urban settings during our in-the-wild deployments; Third, the magical theme enables various social interactions—from solo practice to multiplayer duels—and creates a cohesive narrative across all three MOFA experiences through consistent spell-casting interactions: “The Training” (single-player mode for practicing spell-casting skills), “The Duel” (two-player wizard dueling), and “The Dragon” (three or more players cooperatively hunting a virtual monster).
We adopted an in-the-wild approach (Benford et al., 2013) to maximize ecological validity (Chaytor and Schmitter-Edgecombe, 2003), deploying MOFA across diverse real-world public venues—such as university campuses, tech expos (Hu et al., 2024b, 2023b), urban parks, public building courtyards, and natural wilderness sites—and inviting passersby to participate. We used opportunistic recruitment in these public spaces, recording the spectator-view AR feed of each gameplay session and conducting brief post-play contextual inquiries with both players and onlookers to capture immediate reactions. Additionally, to gain broader perspectives beyond those of on-site participants, we performed a desktop review of online forum discussions generated after sharing our actual gameplay videos. This allowed us to compare firsthand in-the-wild experiences with the more detached insights of remote desktop observers, ultimately examining whether the IMRSP vision, initially sparked by those speculative promotional videos, truly holds up in real-world settings.
@@
Unlike traditional mobile-based AR games, IMRSP requires substantial bodily movement in public spaces. While this enables presence and collaborative excitement, it also introduces distinct ethical, social acceptability, privacy, and safety concerns (Millard et al., 2024; Katins et al., 2024). The online review feedback revealed more concerns than those expressed by on-site participants. As IMRSP games are physically demanding and often challenge established social norms—for example, due to the power imbalance between MRHMD wearers and bystanders (Chung et al., 2023)—the long-term player retention of IMRSP in public spaces remains uncertain. These implications underscore that the path to the ubiquitous IMRSP vision may involve friction unless we address these issues head-on. Our empirical studies provide an initial yet foundational understanding of IMRSP. By examining how MRHMD technologies reshape social norms, dynamics, interactions, and collaborative play in public spaces, our findings inform both future research directions and design guidelines for creating more responsible and inclusive public IMRSP experiences.
@@
  • An initial real-world exploration of social dynamics and phenomena in IMRSP, revealing key benefits and challenges of this novel form of play;
  • A comparative analysis of online reviews versus on-site gameplay experiences, highlighting the friction of social acceptability between idealized visions and practical reality;
  • Design recommendations and a research agenda for future IMRSP development derived from our exploratory work;
  • An open-source framework MOFA for spontaneous co-located multiplayer MR gameplay that facilitates further IMRSP research.
2. Background and Related Work
@@ 2.1. Technology-mediated Street Play
Street play has long been an integral part of urban culture, transforming streets into dynamic spaces where social interactions, creativity, and spontaneity flourish (Tranter and Doyle, 1996). Traditional street play, characterized by informal, unstructured activities like hopscotch, tag, or ball games (Wridt, 2004), utilizes public spaces in ways that challenge the boundaries of social and spatial order within urban environments (Stevens, 2007). Streets become arenas for social interaction, where norms are negotiated and social bonds are strengthened (Whyte, 2010). The spontaneous nature of street play fosters inclusivity and accessibility, enabling people of various ages and backgrounds to participate and interact. Meanwhile, the fields of urban play (Chisik et al., 2022), playable city (Nijholt, 2017b), pervasive play (Ahn et al., 2016), alternate reality games (Kim et al., 2008), and pervasive games (Montola et al., 2009) have all contributed to redefining human-city interactions  (Lee et al., 2021).
@@
Technological advances in network coverage, power consumption, and device miniaturization have paved the way for new forms of outdoor interactive computing (Balestrini et al., 2020). These breakthroughs have spawned diverse deployments and studies of technology use in public spaces. Examples include large interactive displays (Cao et al., 2008; O’Hara et al., 2008), mobile-based mixed reality (Flintham et al., 2003), location-based gaming (such as “Can You See Me Now?” (Benford et al., 2006), geocaching (O’Hara, 2008), and Pokémon GO (Paavilainen et al., 2017)), and human-robot interactions on streets (Pelikan et al., 2024). These technological interventions have transformed street play, introducing innovative forms of engagement that blend physical and digital realms seamlessly (De Souza E Silva, 2009).
The street is a site of socially organized human actions. Studies of behavior in public spaces often draw from Erving Goffman’s work, which identifies social behaviors and the management of impressions in everyday interactions (Goffman, 2021, 1966). Ethnomethodological and conversation analytic studies have examined public settings, emphasizing that social order is fragile and easily disrupted (McGrane, 1994). Shusterman’s “Bodies in the Streets” (Shusterman, 2019) examines how bodily sensations, practices, and perceptions profoundly shape our everyday experiences in urban environments, illuminating the somaesthetic dimensions of city life. These studies highlight how individuals utilize subtle bodily cues in street interactions to approach others, regardless of familiarity (De Stefani and Mondada, 2018). Furthermore, visual information plays a crucial role in how people navigate the street, coordinating movements and avoiding collisions (Hester and Francis, 2003). Together, these insights illustrate the delicate nature of social interactions in urban environments and the embodied practices that facilitate social order amidst potential disruptions.
@@
Informed by such studies, HCI has developed conceptual frameworks for thinking about designing for interaction in public, including research in the wild (Chamberlain et al., 2012), performance-led research (Benford et al., 2013), and design considerations of bystanders and spectators in public interactions (Reeves et al., 2005; Reeves, 2011). These approaches have been instrumental in understanding and shaping user experiences in outdoor settings. Meanwhile, HCI research has documented numerous social phenomena that arise when people interact with technology in public contexts, such as public displays (O’Hara et al., 2008) and public installations (Balestrini et al., 2016). Among these are the honeypot effect (Wouters et al., 2016), where an initial user attracts onlookers and encourages them to engage; showmanship and performance, where users become aware of an audience and amplify their actions  (Benford and Giannachi, 2011; Peltonen et al., 2008; Reeves et al., 2005); social mirroring, in which newcomers imitate observed behaviors (Peltonen et al., 2008). Researchers have also noted social discomfort (Benford et al., 2012), social facilitation (improved performance in front of others) (Peltonen et al., 2008; Miller et al., 2019), social awkwardness (Deterding, 2018; Huggard et al., 2013; Press and Erel, 2023), and inhibition (anxiety or reluctance to participate publicly) (Miller et al., 2019).
2.2. Technology-supported Co-located Bodily Play
@@
In recent years, HCI research has undergone a “somatic turn” (Loke and Schiphorst, 2018). This shift places the human body at the center of interactive technology design, moving beyond traditional user-centered approaches (Dix, 2017) to focus on the body as the primary site of interaction and experience (Höök, 2018). Merleau-Ponty argues that the human body is fundamental to perception and action (Merleau-Ponty, 2008), with proprioception and kinesthesia serving as core movement-related senses (Zhou et al., 2021). Homewood et al. (Homewood et al., 2021) outline three evolving conceptions of the body in HCI: the individual body (embodied interaction), multiple bodies (plurality of experiences), and more-than-human bodies (entangled assemblages). These concepts, particularly entangled assemblages (Forlano, 2017), intercorporeality (Meyer et al., 2017; Stepanova and Reicke, 2024; Patibanda et al., 2024), and post-phenomenology (Wakkary et al., 2018), provide theoretical frameworks for exploring technology-supported co-located bodily play.
@@
The conceptual frameworks of proxemics (Ciolek, 1983) and f-formations (Barua et al., 2021) significantly inform the practical exploration of co-located bodily play. They offer insights into how spatial and social arrangements enhance interaction within shared physical spaces. These principles are instrumental in designing co-located bodily experiences that are socially engaging and spatially harmonious. Research on co-located bodily play (Segura and Isbister, 2015) in exertion games (Kaos et al., 2019) further underscores the importance of multiplayer experiences and social interaction in enhancing game adherence and enjoyment. Co-located bodily play design is rooted in social bodily play (Mueller et al., 2018b, 2019), encompassing three key areas: bodily play (Mueller et al., 2014; Jørgensen and Kaygan, 2024; Martin-Niedecken, 2018; Mueller et al., 2020; Márquez Segura et al., 2013; Isbister et al., 2018), movement-based design (Mueller and Isbister, 2014; van Rheden et al., 2024; Van Delden et al., 2023; Isbister and Mueller, 2015; Garner et al., 2013; Hämäläinen et al., 2015), and social exertion games (Mueller et al., 2017; Sinclair et al., 2007; Márquez Segura et al., 2021; Mueller and Young, 2017; Elvitigala et al., 2024). This body of work collectively emphasizes the integration of physical activity, social interaction, and immersive digital content in creating engaging and harmonious social play experiences.
2.3. Co-located Mixed Reality